### v1

# Asymptotic solutions of generalized Fermat-type equation of signature $(p,p,3)$ over totally real number fields

Abstract.
In this article, we study the asymptotic solutions of the generalized Fermat-type equation of signature $(p,p,3)$ over totally real number fields $K$, i.e., $Ax^{p}+By^{p}=Cz^{3}$ with prime exponent $p$ and $A,B,C\in\mathcal{O}_{K}\setminus\{0\}$. For certain class of fields $K$, we prove that $Ax^{p}+By^{p}=Cz^{3}$ has no asymptotic solutions over $K$ (resp., solutions of certain type over $K$) with restrictions on $A,B,C$ (resp., for all $A,B,C\in\mathcal{O}_{K}\setminus\{0\}$). Finally, we present several local criteria over $K$.
Key words and phrases:
Fermat-Type equations, Signature $(p,p,3)$, Semi-stability, Irreducibility, Modularity, Level lowering
2010 Mathematics Subject Classification:
Primary 11D41, 11R80; Secondary 11F80, 11G05, 11R04
1. Introduction
The study of Diophantine equations is an interesting and extremely fascinating area in number theory. The most important example of Diophantine equation is the Fermat equation $x^{n}+y^{n}=z^{n}$. In 1637, Fermat claimed that the equation $x^{n}+y^{n}=z^{n}$ with positive integers $n\geq 3$ has no non-trivial co-prime integer solutions. This was settled affirmatively by Wiles. The main inputs in the proof were the modularity of semi-stable elliptic curves $E/\mathbb{Q}$ (cf. [Wil95][TW95]), irreducibility of the mod $p$ Galois representations $\bar{\rho}_{E,p}$ attached to $E/\mathbb{Q},p$ (cf. [Maz78]), and the level lowering theorem of Ribet for $\bar{\rho}_{E,p}$ (cf. [Rib90]).
Since then, there has been a lot of progress in understanding the solutions of the generalized Fermat-type equation
$Ax^{p}+By^{q}=Cz^{r}$ (1.1)
with integers $A,B,C$ are co-prime and $p,q,r\geq 2$ are integers with $\frac{1}{p}+\frac{1}{q}+\frac{1}{r}<1$. We call $(p,q,r)$ as the signature of (1.1). In [DG95], Darmon and Granville showed that the generalized Fermat-type equation (1.1), with fixed $A,B,C,p,q,r$, has only finitely many non-trivial co-prime integer solutions.
1.1. Literature Survey:
Throughout this article, $K$ denotes a totally real number field, and $\mathcal{O}_{K}$ denotes the ring of integers of $K$. Let $n\in\mathbb{N}$, and $p\in\mathbb{P}:=\mathrm{Spec}(\mathbb{Z})$. In [DM97], Darmon and Merel proved that the equation $x^{n}+y^{n}=z^{3}$ with exponent $n$ has no non-trivial co-prime integer solutions for $n\geq 3$. In [BVY04], Bennett, Vatsal, and Yazdani studied the integer solution of the generalized Fermat-type equation of signature $(n,n,3)$, i.e., $Ax^{n}+By^{n}=Cz^{3}$, and proved that the equation $x^{n}+y^{n}=pz^{3}$ has no co-prime integer solutions with $|xy|>1$ and prime $n>p^{4p^{2}}$. Recently, in [Moc22], Mocanu studied the asymptotic solutions (of certain type over $K$) for the equation $x^{p}+y^{p}=z^{3}$ with exponent $p$ (cf. Definition 2.4 for the definition of the asymptotic solutions). More recently, in [IKO23], Işik, Kara, and Özman studied the asymptotic solution of certain type over number fields $F$ for the equation $x^{p}+y^{p}=z^{3}$ with exponent $p$, by assuming two modularity conjectures (cf. [IKO23, Conjectures 2.2, 2.3]) when the narrow class number of $F$ is $1$.
In this article, we generalize the work of [Moc22] and study the asymptotic solutions of the generalized Fermat-type equation of signature $(p,p,3)$, i.e., $Ax^{p}+By^{p}=Cz^{3}$ with exponent $p$ over $K$, where $A,B,C\in\mathcal{O}_{K}\setminus\{0\}$. More precisely,
  • In Theorem 2.5, we prove that, for certain class of fields $K$, the equation $Ax^{p}+By^{p}=Cz^{3}$ with prime exponent $p$ has no asymptotic solution in $W_{K}$ (cf. Definition 2.2 for $W_{K}$).
  • In Theorem 3.1, we prove that, for certain class of fields $K$, the equation $Ax^{p}+By^{p}=Cz^{3}$ with prime exponent $p$ has no asymptotic solution in $\mathcal{O}_{K}^{3}\setminus S$ for some explicit set $S\subseteq\mathcal{O}_{K}^{3}$, whenever $3$ is inert in $K$, $v_{\mathfrak{P}}(A)=1$, $v_{\mathfrak{P}}(B)\in\{0,2\}$ for $\mathfrak{P}|3$, and $C\in\mathcal{O}_{K}^{*}$ or $C=uq$ for $u\in\mathcal{O}_{K}^{*}$ and $q\in\mathbb{P}\setminus\{3\}$ (cf. §3.1 for $S$). In fact, $S=\emptyset$ if $A,B,C\in\mathbb{Z}\setminus\{0\}$ (cf. Proposition (3.6)).
The proofs of Theorems 2.53.1 depend upon certain explicit bounds on the solutions of the equation
$\alpha+\beta=\gamma^{3}$ (1.2)
with $\alpha,\beta\in\mathcal{O}_{S_{K}^{\prime}}^{\ast},\gamma\in\mathcal{O}_{S_{K}^{\prime}}$ (cf. §2 for the definitions of $S^{\prime}_{K},\mathcal{O}_{S_{K}^{\prime}},\mathcal{O}_{S_{K}^{\prime}}^{\ast}$). Finally, we provide several local criteria of $K$ for Theorem 2.5.
1.2. Literature for the equations of signature $(p,p,p)$ and $(p,p,2)$
In [FS15a], Freitas and Siksek studied the asymptotic solutions of the Fermat equation $x^{p}+y^{p}=z^{p}$ with exponent $p$ over $K$. In [Dec16], Deconinck extended the work of [FS15a] to the generalized Fermat equation $Ax^{p}+By^{p}=Cz^{p}$ with $A,B,C$ are odd. In [KS24], we studied the asymptotic solutions of the equation $x^{p}+y^{p}=2^{r}z^{p}$ with exponent $p$ and $r\in\mathbb{N}$. In general, there has been a lot of progress, which is summarized below, for the generalized Fermat-type equation of signature $(n,n,2)$ over $K$. Let $r,n\in\mathbb{N}$, $p\in\mathbb{P}:=\mathrm{Spec}(\mathbb{Z})$ and $P:=\mathrm{Spec}(\mathcal{O}_{K})$.
Fermat-type equation $\mathbb{Q}$-solutions asymptotic $K$-solutions
$x^{n}+y^{n}=z^{2}$ [DM97]  [IKO20], [Moc22], [KS24] for $n=p$
$x^{p}+2^{r}y^{p}=z^{2}$ [Ivo03], [Sik03] [KS]
$x^{p}+2^{r}y^{p}=2z^{2}$ [Ivo03] [KS]
$Ax^{n}+By^{n}=Cz^{2}$ [BS04] [KS] for $n=p$ and $C=1$ or $2$
1.3. Our strategy:
In this section, we discuss the modular method, due to Freitas and Siksek [FS15a], to prove Theorems 2.53.1. The proof of Theorem 2.5 is fairly standard. The novelty of our work can be seen in the proof of Theorem 3.1, which is elaborated below.
  1. (1)
    For any non-trivial primitive solution $(a,b,c)\in\mathcal{O}_{K}^{3}$ to the equation $Ax^{p}+By^{p}=Cz^{3}$, we attach a Frey elliptic curve $E:=E_{a,b,c}$ as in (2.4). Then we prove the modularity of $E$ for $p\gg 0$, and $E$ has semi-stable reduction at $\mathfrak{q}\in P\setminus S_{K}^{\prime}$ along with $p|v_{\mathfrak{q}}(\Delta_{E})$. Using [FS15b, Theorem 2], the residual representation $\bar{\rho}_{E,p}$ is irreducible for $p\gg 0$.
  2. (2)
    Let $(a,b,c)\in\mathcal{O}_{K}^{3}\setminus S$. By a result Freitas and Siksek on the image of inertia at $\mathfrak{q}\in P$ (cf. Lemma 2.10) and a result of [Kra90] (cf. Lemma 3.3) for $p\gg 0$, we get that either $p|\#\bar{\rho}_{E,p}(I_{\mathfrak{P}})$ or #$\#\bar{\rho}_{E,p}(I_{\mathfrak{P}})\in\{3,6\}$ for $\mathfrak{P}\in S_{K}$ if $3$ is inert in $K$ and with some assumptions on $A,B,C$ (cf. Lemma 3.4).
  3. (3)
    Now, using the modularity of $E$, irreducibility of $\bar{\rho}_{E,p}$, semi-stable reduction of $E$ away from $S_{K}^{\prime}$ and a level lowering result by  [FS15a, Theorem 7], there exists an elliptic curve $E^{\prime}/K$ having a non-trivial $3$-torsion point with $\bar{\rho}_{E,p}\sim\bar{\rho}_{E^{\prime},p}$ for $p\gg 0$ (cf. Theorem 3.5). Again, using Lemmas 2.10,  3.3, we get either $v_{\mathfrak{P}}(j_{E^{\prime}})<0$ or $v_{\mathfrak{P}}(j_{E^{\prime}})\equiv 0\text{ or }2\pmod{3}$ for $\mathfrak{P}\in S_{K}$. Finally, using a technique of Mocanu in [Moc22], we relate $j_{E^{\prime}}$ in terms of solutions of (1.2), together with (3.2), to get $v_{\mathfrak{P}}(j_{E^{\prime}})\geq 0$ and $v_{\mathfrak{P}}(j_{E^{\prime}})\equiv 1\pmod{3}$ for some $\mathfrak{P}\in S_{K}$ to get a contradiction.
1.4. Preliminaries
Let $P:=\mathrm{Spec}(\mathcal{O}_{K})$, and $\mathfrak{n}$ be an ideal of $\mathcal{O}_{K}$. Let $E/K$ be an elliptic curve of conductor $\mathfrak{n}$. Let $\bar{\rho}_{E,p}:G_{K}:=\mathrm{Gal}(\overline{K}/K)\rightarrow\mathrm{Aut}(E[p])\simeq\mathrm{GL}_{2}(\mathbb{F}_{p})$ be the residual Galois representation of $G_{K}$, induced by the action of $G_{K}$ on the $p$-torsion of $E$. For any Hilbert modular newform $f$ over $K$ of weight $k$, level $\mathfrak{n}$ with coefficient field $\mathbb{Q}_{f}$ and any $\omega\in\mathrm{Spec}(\mathcal{O}_{\mathbb{Q}_{f}})$, let $\bar{\rho}_{f,\omega}:G_{K}\rightarrow\mathrm{GL}_{2}(\mathbb{F}_{\omega})$ be the residual Galois representation attached to $f,\omega$. The following conjecture is a generalization of the Eichler-Shimura theorem over $\mathbb{Q}$.
Conjecture 1.1 (Eichler-Shimura).
Let $f$ be a Hilbert modular newform over $K$ of parallel weight $2$, level $\mathfrak{n}$, and with coefficient field $\mathbb{Q}_{f}=\mathbb{Q}$. Then, there exists an elliptic curve $E_{f}/K$ with conductor $\mathfrak{n}$ having same $L$-function as $f$.
In [Dar04, Theorem 7.7], Darmon proved Conjecture 1.1 over $K$, when $[K:\mathbb{Q}]$ is odd or there exists some $\mathfrak{q}\in P$ such that $v_{\mathfrak{q}}(\mathfrak{n})=1$. In [FS15a, Corollary 2.2], Freitas and Siksek provided a partial answer to Conjecture 1.1 in terms of the residual Galois representations $\bar{\rho}_{E,p}$ attached to $E$. For any $\mathfrak{q}\in P$, let $\Delta_{\mathfrak{q}}$ be the minimal discriminant of $E$ at $\mathfrak{q}$. Let
$\mathfrak{m}_{p}:=\prod_{p|v_{\mathfrak{q}}(\Delta_{\mathfrak{q}}),\ \mathfrak{q}||\mathfrak{n}}\mathfrak{q}\text{ and }\mathfrak{n}_{p}:=\frac{\mathfrak{n}}{\mathfrak{m}_{p}}.$ (1.3)
We end this section with a result of Freitas and Siksek (cf. [FS15a, Theorem 7]) on the level lowering of the residual Galois representations $\bar{\rho}_{E,p}$ over $K$.
Theorem 1.2.
Let $E$ be an elliptic curve over $K$ of conductor $\mathfrak{n}$. Let $p$ be a rational prime. Suppose that the following conditions hold:
  1. (1)
    For $p\geq 5$, the ramification index $e(\mathfrak{q}/p)<p-1$ for all $\mathfrak{q}|p$, and $\mathbb{Q}(\zeta_{p})^{+}\nsubseteq K$;
  2. (2)
    $E/K$ is modular, and $\bar{\rho}_{E,p}$ is irreducible;
  3. (3)
    $E$ is semi-stable at all $\mathfrak{q}|p$, and $p|v_{\mathfrak{q}}(\Delta_{\mathfrak{q}})$ for all $\mathfrak{q}|p$.
Then there exists a Hilbert modular newform $f$ over $K$ of parallel weight $2$, level $\mathfrak{n}_{p}$, and some prime $\omega$ of $\mathbb{Q}_{f}$ such that $\omega|p$ and $\bar{\rho}_{E,p}\sim\bar{\rho}_{f,\omega}$.
2. Solutions of the equation $Ax^{p}+By^{p}=Cz^{3}$ over $W_{K}$
In this section, we study the solutions of the following equation
$Ax^{p}+By^{p}=Cz^{3}$ (2.1)
with prime exponent $p\geq 3$ and $A,B,C\in\mathcal{O}_{K}\setminus\{0\}$. Recall that $P=\mathrm{Spec}(\mathcal{O}_{K})$. Let $S_{K}:=\{\mathfrak{P}\in P:\ \mathfrak{P}|3\}$ and $S_{K}^{\prime}:=\{\mathfrak{P}\in P:\ \mathfrak{P}|3ABC\}$.
Definition 2.1 (Trivial solution).
We call a solution $(a,b,c)\in\mathcal{O}_{K}^{3}$ to the equation (2.1) with exponent $p$ is trivial, if $abc=0$, otherwise non-trivial. We say $(a,b,c)\in\mathcal{O}_{K}^{3}$ is primitive if $a\mathcal{O}_{K}+b\mathcal{O}_{K}+c\mathcal{O}_{K}=\mathcal{O}_{K}$.
Definition 2.2.
Let $W_{K}$ be the set of all non-trivial primitive solutions $(a,b,c)\in\mathcal{O}_{K}^{3}$ to the equation (2.1) with exponent $p$ such that $\mathfrak{P}|ab$ for every $\mathfrak{P}\in S_{K}$.
Remark 2.3.
Let $\mathfrak{P}\in S_{K}$. If $(a,b,c)\in W_{K}$ with $p>v_{\mathfrak{P}}(C)$, then $\mathfrak{P}$ divides exactly one of $a,b$. Otherwise $\mathfrak{P}^{p}|Aa^{p}+Bb^{p}=Cc^{3}$ and since $p>v_{\mathfrak{P}}(C)$, we get $\mathfrak{P}|c$. This is a contradiction to the primitivity of $(a,b,c)$.
2.1. Main result
For any set $S\subseteq P$, let $\mathcal{O}_{S}:=\{\alpha\in K:v_{\mathfrak{P}}(\alpha)\geq 0\text{ for all }\mathfrak{P}\in P\setminus S\}$ be the ring of $S$-integers in $K$ and $\mathcal{O}_{S}^{*}$ be the $S$-units of $\mathcal{O}_{S}$. Let $\mathrm{Cl}_{S}(K):=\mathrm{Cl}(K)/\langle[\mathfrak{P}]\rangle_{\mathfrak{P}\in S}$ and $\mathrm{Cl}_{S}(K)[n]$ be its $n$-torsion points, where $\mathrm{Cl}(K)$ denote the class group of $K$.
Definition 2.4.
We say a Diophantine equation $Ax^{p}+By^{p}=Cz^{3}$ with exponent $p$ has no asymptotic solution in a set $W\subseteq\mathcal{O}_{K}^{3}$, if there exists a constant $V_{K,A,B,C}>0$ (depending on $K,A,B,C$) such that for primes $p>V_{K,A,B,C}$, the equation $Ax^{p}+By^{p}=Cz^{3}$ with exponent $p$ has no non-trivial primitive solution in $W$.
We now show that the equation (2.1) with exponent $p$ has no asymptotic solution in $W_{K}$. More precisely,
Theorem 2.5.
Let $K$ be a totally real field with $\mathrm{Cl}_{S_{K}^{\prime}}(K)[3]=1$. Suppose for every solution $(\alpha,\beta,\gamma)\in\mathcal{O}_{S_{K}^{\prime}}^{\ast}\times\mathcal{O}_{S_{K}^{\prime}}^{\ast}\times\mathcal{O}_{S_{K}^{\prime}}$ to $\alpha+\beta=\gamma^{3}$, there exists $\mathfrak{P}\in S_{K}$ that satisfies
$\left|v_{\mathfrak{P}}\left(\alpha\beta^{-1}\right)\right|\leq 3v_{\mathfrak{P}}(3).$ (2.2)
Then, the equation $Ax^{p}+By^{p}=Cz^{3}$ with exponent $p$ has no asymptotic solution in $W_{K}$.
Remark 2.6.
By [Moc22, Theorem 39], for any finite set $S\subseteq P$, the equation $\alpha+\beta=\gamma^{3}$ with $\alpha,\beta\in\mathcal{O}_{S}^{\ast}$ and $\gamma\in\mathcal{O}_{S}$ has only finitely many solutions. So, the hypothesis (2.2) in Theorem 2.5 needs to be checked only finitely many times.
We say that $S\subseteq P$ is principal if $\mathfrak{P}$ is principal for every $\mathfrak{P}\in S$. If $3$ is inert in $K$, and either $A,B,C\in\mathcal{O}_{K}^{\ast}$ or $A,B,C\in\mathbb{P}$ inert in $K$, then $S_{K}^{\prime}$ is principal. In this case, $\mathrm{Cl}_{S^{\prime}_{K}}(K)=\mathrm{Cl}(K)$, and $\mathrm{Cl}(K)[3]=1$ is equivalent to $3\nmid h_{K}$. The following proposition is a consequence of Theorem 2.5, which will be useful in §4.
Proposition 2.7.
Let $K$ be a field such that $S_{K}^{\prime}=S_{K}$ is principal and $3\nmid h_{K}h_{K(\zeta_{3})}$, where $h_{K},h_{K(\zeta_{3})}$ denote the class numbers of $K$, $K(\zeta_{3})$, respectively. Assume that $3$ is inert or totally ramified in $K$, and $S_{K}=\{\mathfrak{P}\}$. Suppose for every solution $(\alpha,\gamma)\in\mathcal{O}_{S_{K}}^{\ast}\times\mathcal{O}_{S_{K}}$ to $\alpha+1=\gamma^{3}$ with $v_{\mathfrak{P}}(\alpha)\geq 0$ satisfies the inequality
$v_{\mathfrak{P}}(\alpha)\leq 3v_{\mathfrak{P}}(3).$ (2.3)
Then, the equation $Ax^{p}+By^{p}=Cz^{3}$ with exponent $p$ has no asymptotic solution in $W_{K}$.
2.2. Construction of Frey elliptic curves
For any non-trivial and primitive solution $(a,b,c)\in\mathcal{O}_{K}^{3}$ to the equation (2.1) with exponent $p$, the Frey curve $E:=E_{a,b,c}$ is given by
$E:=E_{a,b,c}:Y^{2}+3CcXY+C^{2}Bb^{p}Y=X^{3},$ (2.4)
with $c_{4}=3^{2}C^{3}c(9Aa^{p}+Bb^{p}),\ \Delta_{E}=3^{3}AB^{3}C^{8}(ab^{3})^{p}$ and $j_{E}=3^{3}\frac{Cc^{3}(9Aa^{p}+Bb^{p})^{3}}{AB^{3}(ab^{3})^{p}}$, where $j_{E}$ (resp., $\Delta_{E}$) denote the $j$-invariant (resp., discriminant) of $E$.
2.3. Modularity of Frey elliptic curves
We now use a modularity result of Freitas and Siksek (cf. [FLBS15, Theorem 5]) to prove the modularity of the Frey curve $E:=E_{a,b,c}$ in  (2.4) associated to $(a,b,c)\in W_{K}$ for primes $p\gg 0$.
Theorem 2.8.
Let $(a,b,c)\in W_{K}$ be a solution to the equation (2.1) with exponent $p$ and let $E:=E_{a,b,c}$ be the Frey curve attached to $(a,b,c)$ as in (2.4). Then, there exists a constant $D:=D_{K,A,B,C}$ (depending on $K,A,B,C$) such that for primes $p>D$, $E/K$ is modular.
Proof.
By [FLBS15, Theorem 5], there exist only finitely many elliptic curves over $K$, up to $\bar{K}$-isomorphism, which are not modular. Let $j_{1},\ldots,j_{s}\in K$ be the $j$-invariants of those elliptic curves. Then $j$-invariant of the Frey curve $E$ is given by $j_{E}=3^{3}\frac{Cc^{3}(9Aa^{p}+Bb^{p})^{3}}{AB^{3}(ab^{3})^{p}}=3^{3}\frac{(Aa^{p}+Bb^{p})(9Aa^{p}+Bb^{p})^{3}}{AB^{3}(ab^{3})^{p}}=3^{3}\frac{(1+\mu(E))(9+\mu(E))^{3}}{\mu(E)^{3}}$ for $\mu(E)=\frac{Bb^{p}}{Aa^{p}}$. For each $i=1,2,\ldots,s$, the equation $j_{E}=j_{i}$ has at most four solutions in $K$. So, there exists $\mu_{1},\mu_{2},...,\mu_{n}\in K$ with $n\leq 4s$ such that $E$ is modular for all $\mu(E)\notin\{\mu_{1},\mu_{2},...,\mu_{n}\}$. If $\mu(E)=\mu_{k}$ for some $k\in\{1,2,\ldots,n\}$, then $\left(\frac{b}{a}\right)^{p}=\frac{A\mu_{k}}{B}$. This equation determines $p$ uniquely, denote it by $p_{k}$. Suppose $q\neq l$ are primes such that $\left(\frac{b}{a}\right)^{q}=\left(\frac{b}{a}\right)^{l}$, which means $\left(\frac{b}{a}\right)$ is a root of unity. Since $K$ is totally real, we get $b=\pm a$. By Remark 2.3, this cannot happen if we choose $p>v_{\mathfrak{P}}(C)$ for $\mathfrak{P}\in S_{K}$. Now, the proof of the theorem follows by taking $D=\max\big{\{}p_{1},...,p_{m},v_{\mathfrak{P}}(C)\big{\}}$. ∎
2.4. Reduction type of Frey elliptic curves
The following lemma characterizes the type of reduction of the Frey curve $E:=E_{a,b,c}$ at primes $\mathfrak{q}$ away from $S_{K}^{\prime}$.
Lemma 2.9.
Let $(a,b,c)\in\mathcal{O}_{K}^{3}$ be a non-trivial primitive solution to the equation (2.1) with exponent $p$, and let $E$ be the associated Frey curve. Then at all primes $\mathfrak{q}$ away from $S_{K}^{\prime}$, $E$ is minimal, semi-stable at $\mathfrak{q}$ and satisfies $p|v_{\mathfrak{q}}(\Delta_{E})$. Let $\mathfrak{n}$ be the conductor of $E$, and $\mathfrak{n}_{p}$ be as in (1.3). Then,
$\mathfrak{n}=\prod_{\mathfrak{P}\in S_{K}^{\prime}}\mathfrak{P}^{r_{\mathfrak{P}}}\prod_{\mathfrak{q}|ab,\ \mathfrak{q}\notin S_{K}^{\prime}}\mathfrak{q},\ \mathfrak{n}_{p}=\prod_{\mathfrak{P}\in S_{K}^{\prime}}\mathfrak{P}^{r_{\mathfrak{P}}^{\prime}},$ (2.5)
where $0\leq r_{\mathfrak{P}}^{\prime}\leq r_{\mathfrak{P}}$ with $r_{\mathfrak{P}}\leq 2+3v_{\mathfrak{P}}(3)$ for $\mathfrak{P}|3$ and $r_{\mathfrak{P}}\leq 2+6v_{\mathfrak{P}}(2)$ for $\mathfrak{P}\nmid 3$.
Proof.
Let $\mathfrak{q}\in P\setminus S_{K}^{\prime}$.
  • If $\mathfrak{q}\not|\Delta_{E}$, then $E$ has good reduction at $\mathfrak{q}$ and $p|v_{\mathfrak{q}}(\Delta_{E})=0$.
  • If $\mathfrak{q}|\Delta_{E}=3^{3}AB^{3}C^{8}(ab^{3})^{p}$, then $\mathfrak{q}$ exactly divides one of $a$ and $b$, since $(a,b,c)$ is primitive and $\mathfrak{q}\nmid 3ABC$. This implies $\mathfrak{q}\nmid c$, hence $\mathfrak{q}\nmid c_{4}=3^{2}C^{3}c(9Aa^{p}+Bb^{p})$. Therefore, $E$ is minimal, and $E$ has multiplicative reduction at $\mathfrak{q}$.
Since $v_{\mathfrak{q}}(\Delta_{E})=pv_{\mathfrak{q}}(ab^{3})$, $p|v_{\mathfrak{q}}(\Delta_{E})$. By the definition of $\mathfrak{n}_{p}$ in (1.3), we get $\mathfrak{q}\nmid\mathfrak{n}_{p}$ for all $\mathfrak{q}\notin S_{K}^{\prime}$. Finally, for $\mathfrak{P}\in S_{K}^{\prime}$, the bounds on $r_{\mathfrak{P}}$ follow from [Sil94, Theorem IV.10.4]. ∎
2.4.1. Type of reduction with image of inertia
Now, we recall  [FS15a, Lemma 3.4], which will be useful for the types of the reduction of the Frey curve at $\mathfrak{q}\in P$.
Lemma 2.10.
Let $E/K$ be an elliptic curve and $p>5$ be a prime. For $\mathfrak{q}\in P$ with $\mathfrak{q}\nmid p$, $E$ has potentially multiplicative reduction at $\mathfrak{q}$ and $p\nmid v_{\mathfrak{q}}(j_{E})$ if and only if $p|\#\bar{\rho}_{E,p}(I_{\mathfrak{q}})$.
The following lemma determines the type of reduction of the Frey curve $E$ at primes $\mathfrak{q}\nmid 3pABC$.
Lemma 2.11.
Let $(a,b,c)\in\mathcal{O}_{K}^{3}$ be a non-trivial primitive solution to the equation (2.1) with exponent $p>5$, and let $E$ be the associated Frey curve. Suppose $\mathfrak{q}\in P$ with $\mathfrak{q}\nmid 3pABC$. Then $p\nmid\#\bar{\rho}_{E,p}(I_{\mathfrak{q}})$.
Proof.
By Lemma 2.10, it is enough to show that either $v_{\mathfrak{q}}(j_{E})\geq 0$ or $p|v_{\mathfrak{q}}(j_{E})$. Recall that $\Delta_{E}=3^{3}AB^{3}C^{8}(ab^{3})^{p}$ and $c_{4}=3^{2}C^{3}c(9Aa^{p}+Bb^{p})$.
  • If $\mathfrak{q}\nmid\Delta_{E}$, then $E$ has good reduction at $\mathfrak{q}$, and hence $v_{\mathfrak{q}}(j_{E})\geq 0$.
  • If $\mathfrak{q}|\Delta_{E}$ then $\mathfrak{q}|ab$, and hence $\mathfrak{q}$ divides exactly one of $a$ and $b$. Therefore, $\mathfrak{q}\nmid c_{4}$ and $p|v_{\mathfrak{q}}(j_{E})=-pv_{\mathfrak{q}}(ab^{3})$.
Hence, we are done with the proof of the lemma. ∎
We will now discuss the type of reduction of $E_{a,b,c}$ at $\mathfrak{P}\in S_{K}$ with $(a,b,c)\in W_{K}$.
Lemma 2.12.
Let $\mathfrak{P}\in S_{K}$. Let $(a,b,c)\in W_{K}$ be a solution to (2.1) with exponent $p>\max\left\{3v_{\mathfrak{P}}(3)+v_{\mathfrak{P}}(ABC),|3v_{\mathfrak{P}}(3)+v_{\mathfrak{P}}(BA^{-1})|,|4v_{\mathfrak{P}}(3)+v_{\mathfrak{P}}(AB^{-1})|\right\}.$ Let $E:=E_{a,b,c}$ be the associated Frey curve. Then $\ v_{\mathfrak{P}}(j_{E})<0$ and $p\nmid v_{\mathfrak{P}}(j_{E})$, equivalently $p|\#\bar{\rho}_{E,p}(I_{\mathfrak{P}})$.
Proof.
Recall that $j_{E}=3^{3}\frac{Cc^{3}(9Aa^{p}+Bb^{p})^{3}}{AB^{3}(ab^{3})^{p}}=3^{3}\frac{(Aa^{p}+Bb^{p})(9Aa^{p}+Bb^{p})^{3}}{AB^{3}(ab^{3})^{p}}.$ Since $p>v_{\mathfrak{P}}(C)$, by Remark 2.3, we have $\mathfrak{P}|a$ or $\mathfrak{P}|b$ but not both.
  • If $\mathfrak{P}|a$, then $\mathfrak{P}\nmid b$. Since $p>v_{\mathfrak{P}}(B)$, $v_{\mathfrak{P}}(j_{E})=3v_{\mathfrak{P}}(3)+v_{\mathfrak{P}}(B)+3v_{\mathfrak{P}}(B)-v_{\mathfrak{P}}(AB^{3})-pv_{\mathfrak{P}}(a)=3v_{\mathfrak{P}}(3)+v_{\mathfrak{P}}(BA^{-1})-pv_{\mathfrak{P}}(a)$. Since $p>|3v_{\mathfrak{P}}(3)+v_{\mathfrak{P}}(BA^{-1})|$, $v_{\mathfrak{P}}(j_{E})<0$ and $p\nmid v_{\mathfrak{P}}(j_{E})$.
  • If $\mathfrak{P}|b$, then $\mathfrak{P}\nmid a$. Since $p>3v_{\mathfrak{P}}(3)+v_{\mathfrak{P}}(A)$, $v_{\mathfrak{P}}(j_{E})=3v_{\mathfrak{P}}(3)+v_{\mathfrak{P}}(A)+9v_{\mathfrak{P}}(3)+3v_{\mathfrak{P}}(A)-v_{\mathfrak{P}}(AB^{3})-3pv_{\mathfrak{P}}(b)=3\left(4v_{\mathfrak{P}}(3)+v_{\mathfrak{P}}(AB^{-1})-pv_{\mathfrak{P}}(b)\right)$. Since $p>|4v_{\mathfrak{P}}(3)+v_{\mathfrak{P}}(AB^{-1})|$, $v_{\mathfrak{P}}(j_{E})<0$ and $p\nmid v_{\mathfrak{P}}(j_{E})$.
Hence, by Lemma 2.10, we get $p|\#\bar{\rho}_{E,p}(I_{\mathfrak{P}})$. ∎
2.5. Proof of Theorem 2.5.
The proof of this theorem depends on the following auxiliary result.
Theorem 2.13.
Let $K$ be a totally real field. Then, there is a constant $V=V_{K,A,B,C}>0$ (depending on $K,A,B,C$) such that the following hold. Let $(a,b,c)\in W_{K}$ be a solution to the equation (2.1) with exponent $p>V$, and let $E$ be the Frey curve as in (2.4). Then, there exists an elliptic curve $E^{\prime}/K$ such that:
  1. (1)
    $E^{\prime}/K$ has good reduction away from $S_{K}^{\prime}$ and has a non-trivial $3$-torsion point;
  2. (2)
    $\bar{\rho}_{E,p}\sim\bar{\rho}_{E^{\prime},p}$, and $v_{\mathfrak{P}}(j_{E^{\prime}})<0$ for $\mathfrak{P}\in S_{K}$.
Proof of Theorem 2.13.
By Theorem 2.8, $E$ is modular for primes $p>D:=D_{K,A,B,C}$ with $D\gg 0$. By Lemma 2.9, $E$ is semi-stable away from $S_{K}^{\prime}$. If necessary, we can take the Galois closure of $K$ to ensure that $\bar{\rho}_{E,p}$ is irreducible for $p\gg 0$ (cf. [FS15b, Theorem 2]). By Theorem 1.2, there exist a Hilbert modular newform $f$ of parallel weight $2$, level $\mathfrak{n}_{p}$ and some prime $\omega$ of $\mathbb{Q}_{f}$ such that $\omega|p$ and $\bar{\rho}_{E,p}\sim\bar{\rho}_{f,\omega}$ for $p\gg 0$. By allowing $p$ to be sufficiently large, we can assume $\mathbb{Q}_{f}=\mathbb{Q}$. This step uses standard ideas originally due to Mazur that can be found in [BS04, §4], [Coh07, Proposition 15.4.2], and [FS15a, §4]. Let $\mathfrak{P}\in S_{K}$. Then $E$ has potential multiplicative reduction at $\mathfrak{P}$ and $p|\#\bar{\rho}_{E,p}(I_{\mathfrak{P}})$ for $p\gg 0$ (cf. Lemma 2.12). The existence of $E_{f}$ then follows from [FS15a, Corollary 2.2] for all $p\gg 0$ after leaving primes $p$ with $p\mid\left(\text{Norm}(K/\mathbb{Q})(\mathfrak{P})\pm 1\right)$. Therefore, $\bar{\rho}_{E,p}\sim\bar{\rho}_{E_{f},p}$ for some elliptic curve $E_{f}$ with conductor $\mathfrak{n}_{p}$ for $p>V=V_{K,A,B,C}$, where $V_{K,A,B,C}$ is the maximum of all the above implicit/explicit lower bounds.
  • Since the conductor of $E_{f}$ is $\mathfrak{n}_{p}$ given in (2.5), $E_{f}$ has good reduction away from $S_{K}^{\prime}$. Now, arguing as in [Moc22, page 1247], we can enlarge the constant $V$ and by possibly replacing $E_{f}$ with an isogenous curve, say $E^{\prime}$, we get $E^{\prime}/K$ has a non-trivial $3$-torsion point. Since $E_{f}\sim E^{\prime}$, $E^{\prime}$ has good reduction away from $S_{K}^{\prime}$.
  • Since $E_{f}$ is isogenous to $E^{\prime}$ and $\bar{\rho}_{E,p}\sim\bar{\rho}_{E_{f},p}$ implies $\bar{\rho}_{E,p}\sim\bar{\rho}_{E^{\prime},p}$. As a result, we obtain $p|\#\bar{\rho}_{E,p}(I_{\mathfrak{P}})=\#\bar{\rho}_{E^{\prime},p}(I_{\mathfrak{P}})$ for any $\mathfrak{P}\in S_{K}$. Finally, by Lemma 2.10, we have $v_{\mathfrak{P}}(j_{E^{\prime}})<0$ for any $\mathfrak{P}\in S_{K}$.
This completes the proof of the theorem. ∎
We now prove Theorem 2.5, and its inspired from that of [Moc22, Theorem 3].
Proof of Theorem 2.5.
Suppose $(a,b,c)\in W_{K}$ is a solution to the equation (2.1) with exponent $p>V$, where $V=V_{K,A,B,C}$ be the constant as in Theorem 2.13. By Theorem 2.13, there exists an elliptic curve $E^{\prime}/K$ having a non-trivial $3$-torsion point and good reduction away from $S_{K}^{\prime}$. Then the elliptic curve $E^{\prime}/K$ has a model of the form
$E^{\prime}:y^{2}+a^{\prime}xy+b^{\prime}y=x^{3}+cx^{2}+dx$ (2.6)
for some $a^{\prime},b^{\prime}\in K$ with $j$-invariant $j_{E^{\prime}}=2^{8}\frac{a^{\prime 3}(a^{\prime 3}-24b^{\prime})^{3}}{b^{\prime 3}(a^{\prime 3}-27b^{\prime})}$. Since $E^{\prime}$ has good reduction away from $S_{K}^{\prime}$, $j_{E^{\prime}}\in\mathcal{O}_{S_{K}^{\prime}}$. Take $\lambda:=\frac{a^{\prime 3}}{b^{\prime}}$ and $\mu:=\lambda-27$. Then $\lambda\in\mathcal{O}_{S_{K}^{\prime}}$ and $\mu\in\mathcal{O}_{S_{K}^{\prime}}^{\ast}$ (cf. [Moc22, Lemma 16(ii)]). By [Moc22, Lemma-17(ii)], we get $\lambda\mathcal{O}_{K}=I^{3}J$ for some fractional ideal $I$ and $S_{K}^{\prime}$-ideal $J$. Since $J$ is $S_{K}^{\prime}$-ideal, $1=[I]^{3}\in\mathrm{Cl}_{S_{K}^{\prime}}(K)$. By hypothesis $\mathrm{Cl}_{S_{K}^{\prime}}(K)[3]=1$ which gives $I=\gamma I_{1}$ for some $\gamma\in\mathcal{O}_{K}$ and $S_{K}^{\prime}$-ideal $I_{1}$. Thus, $\lambda\mathcal{O}_{K}=\gamma^{3}I_{1}^{3}J$ and hence $(\frac{\lambda}{\gamma^{3}})\mathcal{O}_{K}$ is an $S_{K}^{\prime}$-ideal. Therefore, $u=\frac{\lambda}{\gamma^{3}}\in\mathcal{O}_{S_{K}^{\prime}}^{\ast}$. Now, divide the equation $\mu+27=\lambda$ by $u$ to obtain $\alpha+\beta=\gamma^{3}$, where $\alpha=\frac{\mu}{u}\in\mathcal{O}_{S_{K}^{\prime}}^{\ast}$ and $\beta=\frac{27}{u}\in\mathcal{O}_{S_{K}^{\prime}}^{\ast}$, which implies $\alpha\beta^{-1}=\frac{\mu}{27}$. By (2.2), there exists $\mathfrak{P}\in S_{K}$ with $|v_{\mathfrak{P}}(\alpha\beta^{-1})|=|v_{\mathfrak{P}}(\frac{\mu}{27})|\leq 3v_{\mathfrak{P}}(3)$. This means
$0\leq v_{\mathfrak{P}}(\mu)\leq 6v_{\mathfrak{P}}(3).$ (2.7)
We now show that the bounds on $v_{\mathfrak{P}}(\mu)$ would imply that $v_{\mathfrak{P}}(j_{E^{\prime}})\geq 0$. Write $j_{E^{\prime}}$ in terms of $\mu$ yields $j_{E^{\prime}}=\frac{(\mu+27)(\mu+3)^{3}}{\mu}$, which means
$v_{\mathfrak{P}}(j_{E^{\prime}})=v_{\mathfrak{P}}(\mu+27)+3v_{\mathfrak{P}}(\mu+3)-v_{\mathfrak{P}}(\mu).$ (2.8)
  • If $0\leq v_{\mathfrak{P}}(\mu)\leq v_{\mathfrak{P}}(3)$, then $v_{\mathfrak{P}}(\mu+27)=v_{\mathfrak{P}}(\mu)$ and $v_{\mathfrak{P}}(\mu+3)\geq v_{\mathfrak{P}}(\mu)$. By (2.8), we get $v_{\mathfrak{P}}(j_{E^{\prime}})=3v_{\mathfrak{P}}(\mu)\geq 0$.
  • If $v_{\mathfrak{P}}(3)<v_{\mathfrak{P}}(\mu)\leq 3v_{\mathfrak{P}}(3)$, then $v_{\mathfrak{P}}(\mu+27)\geq v_{\mathfrak{P}}(\mu)$ and $v_{\mathfrak{P}}(\mu+3)=v_{\mathfrak{P}}(3)$. By (2.8), we have $v_{\mathfrak{P}}(j_{E^{\prime}})>0$.
  • If $3v_{\mathfrak{P}}(3)<v_{\mathfrak{P}}(\mu)\leq 6v_{\mathfrak{P}}(3)$, then $v_{\mathfrak{P}}(\mu+27)=3v_{\mathfrak{P}}(3)$ and $v_{\mathfrak{P}}(\mu+3)=v_{\mathfrak{P}}(3)$. By (2.8), we have $v_{\mathfrak{P}}(j_{E^{\prime}})=6v_{\mathfrak{P}}(3)-v_{\mathfrak{P}}(\mu)\geq 0$.
In all cases, we get $v_{\mathfrak{P}}(j_{E^{\prime}})\geq 0$, which is a contradiction to Theorem 2.13. This completes the proof of the theorem. ∎
Now, we are in a position to prove Proposition 2.7, and its proof is inspired from that of [Moc22, Theorem 11].
Proof of Proposition 2.7.
Let $S_{K}=\{\mathfrak{P}\}$. By Theorem 2.5, it suffices to show that for every solution $(\alpha,\beta,\gamma)\in\mathcal{O}_{S_{K}}^{\ast}\times\mathcal{O}_{S_{K}}^{\ast}\times\mathcal{O}_{S_{K}}$ to the equation $\alpha+\beta=\gamma^{3}$ satisfies $|v_{\mathfrak{P}}(\alpha\beta^{-1})|\leq 3v_{\mathfrak{P}}(3)$. If necessary, by scaling cubic powers of $\mathfrak{P}$ and swapping $\alpha,\beta$, we can assume $0\leq v_{\mathfrak{P}}(\beta)\leq v_{\mathfrak{P}}(\alpha)$ with $v_{\mathfrak{P}}(\beta)=0$ or $1$ or $2$.
  1. (1)
    Suppose $v_{\mathfrak{P}}(\beta)=1$ or $2$. If $v_{\mathfrak{P}}(\alpha)>v_{\mathfrak{P}}(\beta)$, then $v_{\mathfrak{P}}(\gamma^{3})=v_{\mathfrak{P}}(\alpha+\beta)=v_{\mathfrak{P}}(\beta)$, which cannot happen since $v_{\mathfrak{P}}(\gamma^{3})$ is multiple of $3$. So, $v_{\mathfrak{P}}(\alpha)=v_{\mathfrak{P}}(\beta)$. Hence, $\left|v_{\mathfrak{P}}(\alpha\beta^{-1})\right|=0<3v_{\mathfrak{P}}(3)$.
  2. (2)
    Suppose $v_{\mathfrak{P}}(\beta)=0$. Then $\beta\in\mathcal{O}_{K}^{\ast}$.
    • If $\beta$ is a cube, then divide the equation $\alpha+\beta=\gamma^{3}$ by $\beta$ to obtain an equation of the form $\alpha^{\prime}+1=\gamma^{\prime 3}$, where $\alpha^{\prime}=\alpha\beta^{-1}\in\mathcal{O}_{S_{K}^{\prime}}^{\ast}$ with $v_{\mathfrak{P}}(\alpha^{\prime})\geq 0$, and $\gamma^{\prime}\in\mathcal{O}_{S_{K}^{\prime}}$. By (2.3), we obtain $|v_{\mathfrak{P}}(\alpha\beta^{-1})|\leq 3v_{\mathfrak{P}}(3)$.
    • Suppose $\beta$ is not a cube. If $v_{\mathfrak{P}}(\alpha)\leq 3v_{\mathfrak{P}}(3)$, then we are done. Otherwise, $v_{\mathfrak{P}}(\alpha)>3v_{\mathfrak{P}}(3)>1$. This gives $\alpha\equiv 0\pmod{3^{3}}$ and $\gamma^{3}=\alpha+\beta\equiv\beta\pmod{3^{3}}$. Since $v_{\mathfrak{P}}(\gamma^{3})=v_{\mathfrak{P}}(\alpha+\beta)=0$ and $S_{K}=S_{K}^{\prime}$, we get $\gamma\in\mathcal{O}_{K}$. The field $L=K(\zeta_{3},\beta^{\frac{1}{3}})$ is a degree $3$ extension of $K(\zeta_{3})$. We will now show that $L$ is unramified at $3$ to get a contradiction to $3\nmid h_{K(\zeta_{3})}$. Consider an element $\theta:=\frac{\gamma^{2}+\gamma\zeta_{3}\beta^{\frac{1}{3}}+\zeta_{3}^{2}\beta^{\frac{1}{3}}}{3}$. The minimal polynomial of $\theta$ is $m_{\theta}(x)=x^{3}+\frac{\gamma(\gamma^{3}-\beta)}{3}x^{2}-\gamma^{2}x-\frac{(\gamma^{3}-\beta)^{2}}{27}$. Then $m_{\theta}(x)\in\mathcal{O}_{K}[x]$ with discriminant $\Delta_{\theta}=-\frac{2\gamma^{3}(\gamma^{3}-\beta)^{3}}{3^{5}}-\frac{4\gamma^{3}(\gamma^{3}-\beta)^{5}}{3^{9}}+\frac{\gamma^{6}(\gamma^{3}-\beta)^{2}}{3^{2}}-4\gamma^{6}-\frac{(\gamma^{3}-\beta)^{4}}{3^{3}}$. Since $\Delta_{\theta}\equiv-4\gamma^{6}\pmod{3}$ and $v_{\mathfrak{P}}(\gamma^{3})=0$, $L$ is unramified at $3$, which contradicts our hypothesis that $3\nmid h_{K(\zeta_{3})}$.
This completes the proof of the proposition. ∎
3. Solutions of $Ax^{p}+By^{p}=Cz^{3}$ over $K$
In this section, we shall examine the $K$-solutions of the equation
$Ax^{p}+By^{p}=Cz^{3}$ (3.1)
with exponent $p$, where $A,B,C\in\mathcal{O}_{K}\setminus\{0\}$. Throughout this section, we assume $C\in\mathcal{O}_{K}^{*}$ or $C=uq$ with $u\in\mathcal{O}_{K}^{\ast}$ and $q\in\mathbb{P}\setminus\{3\}$. In both cases, $v_{\mathfrak{P}}(C)=0$. Recall, $S_{K}^{\prime}=\{\mathfrak{P}\in P:\mathfrak{P}|3ABC\}$. Let $h_{K}^{+}$ be the narrow class number of $K$.
3.1. Main result
We write $(ES)$ for “either $[K:\mathbb{Q}]\equiv 1\pmod{2}$ or Conjecture 1.1 holds for $K$”. Let $S\subseteq\mathcal{O}^{3}_{K}$ be the set of all solutions of (3.1) of the form $\left\{(u,\pm u,c):u\in\mathcal{O}_{K}^{*},\ c\in\mathcal{O}_{K}\setminus\{0\}\right\}.$ We now show that the equation (3.1) with exponent $p$ has no asymptotic solution in $\mathcal{O}_{K}^{3}\setminus S$. More precisely,
Theorem 3.1.
Let $K$ be a totally real field satisfying $(ES)$ with $\mathrm{Cl}_{S_{K}^{\prime}}(K)[3]=1$. Assume $3$ is inert in $K$ and let $S_{K}=\{\mathfrak{P}\}$. Suppose every solution $(\alpha,\beta,\gamma)\in\mathcal{O}_{S_{K}^{\prime}}^{\ast}\times\mathcal{O}_{S_{K}^{\prime}}^{\ast}\times\mathcal{O}_{S_{K}^{\prime}}$ to $\alpha+\beta=\gamma^{3}$ satisfies
$v_{\mathfrak{P}}(\alpha\beta^{-1})=2.$ (3.2)
Further, if $v_{\mathfrak{P}}(A)=1$ and $v_{\mathfrak{P}}(B)\in\{0,2\}$, then the equation $Ax^{p}+By^{p}=Cz^{3}$ with exponent $p$ has no asymptotic solution in $\mathcal{O}_{K}^{3}\setminus S$.
3.2. Modularity of of Frey elliptic curves
We now prove the modularity of the Frey curve $E:=E_{a,b,c}$ in (2.4) associated to any non-trivial primitive solution $(a,b,c)\in$ $\mathcal{O}_{K}^{3}\setminus S$ for primes $p\gg 0$.
Theorem 3.2.
Let $(a,b,c)\in\mathcal{O}_{K}^{3}\setminus S$ be a non-trivial primitive solution to the equation (3.1) with exponent $p$, and let $E:=E_{a,b,c}$ be the associated Frey curve. Then there exists a constant $D:=D_{K,A,B,C}$ (depending on $K,A,B,C$) such that for primes $p>D$, $E/K$ is modular.
Proof.
Arguing as in the proof of Theorem 2.8, there exists $\mu_{k}\in K$ with $1\leq k\leq n$ such that $E/K$ is modular for all $\mu(E)\notin\{\mu_{1},\mu_{2},...,\mu_{n}\}$. If $\mu=\mu_{k}$ for some $k\in\{1,2,\ldots,n\}$, then $\left(\frac{b}{a}\right)^{p}=\frac{A\mu_{k}}{B}$. The above equation determines $p$ uniquely, denote it $p_{k}$. Otherwise, we get $b=\pm a$. Now, we will show that $a=u,\ b=\pm u$ for some $u\in\mathcal{O}_{K}^{\ast}.$ Let $\mathfrak{q}\in P$. If $\mathfrak{q}|a$, then $\mathfrak{q}^{p}|Aa^{p}+Bb^{p}=Cc^{3}$. If $C\in\mathcal{O}_{K}^{\ast}$, then $\mathfrak{q}|c$. If $C=uq$ with $u\in\mathcal{O}_{K}^{\ast}$ and $q\in\mathbb{P}$, then $\mathfrak{q}^{p}|qc^{3}$. Taking $p>[K:\mathbb{Q}]$, we have $\mathfrak{q}|c$. In both cases, we get a contradiction to the primitivity of $(a,b,c)$. Therefore, $a\in\mathcal{O}_{K}^{\ast}$. Hence $(a,b,c)\in S$, which is a contradiction. Arguing as in the proof of Theorem 2.8, the proof of the theorem follows by taking $D=\max\big{\{}p_{1},...,p_{m},[K:\mathbb{Q}]\big{\}}$. ∎
3.3. Reduction type of Frey elliptic curves
The following lemma will be useful for the reduction of the Frey curve $E_{a,b,c}$ with $(a,b,c)\in$ $\mathcal{O}_{K}^{3}$ at $\mathfrak{P}\in S_{K}$.
Lemma 3.3.
Let $E/K$ be an elliptic curve and $p\geq 5$ be a prime. Assume $3$ is unramified in $K$. Suppose $E$ has potential good reduction at $\mathfrak{P}$ for some $\mathfrak{P}\in S_{K}$.
  1. (1)
    If $v_{\mathfrak{P}}(\Delta_{E})=4$ or $10$, then #$\#\bar{\rho}_{E,p}(I_{\mathfrak{P}})=3$ or $6$.
  2. (2)
    If #$\#\bar{\rho}_{E,p}(I_{\mathfrak{P}})=3$ or $6$, then $v_{\mathfrak{P}}(\Delta_{E})\in\{4,6,10,12\}$.
Proof.
This lemma is a special case of [Kra90, Corollaire to Théoréme 1]. ∎
The following lemma specifies the type of reduction of the Frey curve $E:=E_{a,b,c}$ given in  (2.4) at $\mathfrak{P}\in S_{K}$ when $(a,b,c)\in$ $\mathcal{O}_{K}^{3}$. More precisely,
Lemma 3.4.
Assume $3$ is inert in $K$ and let $S_{K}=\{\mathfrak{P}\}$. Let $(a,b,c)\in\mathcal{O}_{K}^{3}$ be a non-trivial primitive solution to the equation (3.1) with exponent
$p>\max\left\{3v_{\mathfrak{P}}(3)+v_{\mathfrak{P}}(AB),|3v_{\mathfrak{P}}(3)+v_{\mathfrak{P}}(BA^{-1})|,|4v_{\mathfrak{P}}(3)+v_{\mathfrak{P}}(AB^{-1})|\right\}.$
Let $E$ be the associated Frey curve. If $v_{\mathfrak{P}}(A)=1$ and $v_{\mathfrak{P}}(B)\in\{0,2\}$, then either $p|\#\bar{\rho}_{E,p}(I_{\mathfrak{P}})$ or #$\#\bar{\rho}_{E,p}(I_{\mathfrak{P}})\in\{3,6\}$.
Proof.
If $\mathfrak{P}|ab$, then by Lemma 2.12, we have $p|\#\bar{\rho}_{E,p}(I_{\mathfrak{P}})$. Suppose $\mathfrak{P}\nmid ab$. Now, recall that $j_{E}=3^{3}\frac{Cc^{3}(9Aa^{p}+Bb^{p})^{3}}{AB^{3}(ab^{3})^{p}}=3^{3}\frac{(Aa^{p}+Bb^{p})(9Aa^{p}+Bb^{p})^{3}}{AB^{3}(ab^{3})^{p}}.$ Then $v_{\mathfrak{P}}(j_{E})=3+v_{\mathfrak{P}}(Aa^{p}+Bb^{p})+3v_{\mathfrak{P}}(9Aa^{p}+Bb^{p})-v_{\mathfrak{P}}(AB^{3})$. Now,
  • if $v_{\mathfrak{P}}(B)=0$, then $v_{\mathfrak{P}}(j_{E})=3-v_{\mathfrak{P}}(A)>0$,
  • if $v_{\mathfrak{P}}(B)=2$, then $v_{\mathfrak{P}}(j_{E})\geq 3+v_{\mathfrak{P}}(A)+3v_{\mathfrak{P}}(B)-v_{\mathfrak{P}}(AB^{3})>0$. Hence, $E$ has potential good reduction at $\mathfrak{P}$.
Since $\Delta_{E}=3^{3}AB^{3}C^{8}(ab^{3})^{p}$, $v_{\mathfrak{P}}(\Delta_{E})=3+v_{\mathfrak{P}}(AB^{3})$. The hypothesis on the valuations of $A,B$ implies $v_{\mathfrak{P}}(AB^{3})=1$ or $7$, hence $v_{\mathfrak{P}}(\Delta_{E})=4$ or $10$. By Lemma 3.3, we have #$\#\bar{\rho}_{E,p}(I_{\mathfrak{P}})=3$ or $6$. So, we are done with the proof of the lemma. ∎
3.4. Proof of Theorem 3.1
The proof of this theorem depends on the following result.
Theorem 3.5.
Let $K$ be a totally real field satisfying $(ES)$. Assume $3$ is inert in $K$, and $v_{\mathfrak{P}}(A)=1$, $v_{\mathfrak{P}}(B)\in\{0,2\}$. Then, there is a constant $V=V_{K,A,B,C}>0$ (depending on $K,A,B,C$) such that the following hold. Let $(a,b,c)\in\mathcal{O}_{K}^{3}\setminus S$ be a non-trivial primitive solution to the equation (3.1) with exponent $p>V$, and let $E$ be the Frey curve as in  (2.4). Then there exists an elliptic curve $E^{\prime}/K$ such that:
  1. (1)
    $E^{\prime}/K$ has good reduction away from $S_{K}^{\prime}$ and has a non-trivial $3$-torsion point, and $\bar{\rho}_{E,p}\sim\bar{\rho}_{E^{\prime},p}$;
  2. (2)
    For $\mathfrak{P}\in S_{K}$, either $v_{\mathfrak{P}}(j_{E^{\prime}})<0$ or $v_{\mathfrak{P}}(j_{E^{\prime}})\equiv 0\text{ or }2\pmod{3}$.
Proof.
Arguing as in the proof of Theorem 2.13, the first part of Theorem 3.5 follows from  [FS15b, Theorem 2], Theorem 3.2, Lemma 2.9 and Theorem 1.2. Let $\mathfrak{P}\in S_{K}$ be the unique prime lying above $3$. If $p|\#\bar{\rho}_{E,p}(I_{\mathfrak{P}})=\#\bar{\rho}_{E^{\prime},p}(I_{\mathfrak{P}})$, then by Lemma 2.10, we get $v_{\mathfrak{P}}(j_{E^{\prime}})<0$. If $p\nmid\#\bar{\rho}_{E,p}(I_{\mathfrak{P}})$, then by Lemma 3.4, we conclude that #$\#\bar{\rho}_{E^{\prime},p}(I_{\mathfrak{P}})=\#\bar{\rho}_{E,p}(I_{\mathfrak{P}})\in\{3,6\}$. If $v_{\mathfrak{P}}(j_{E^{\prime}})<0$, then we are done. If $v_{\mathfrak{P}}(j_{E^{\prime}})\geq 0$, then by Lemma 3.3, we have $v_{\mathfrak{P}}(\Delta_{E^{\prime}})\equiv 0\text{ or }1\pmod{3}$. Since $j_{E^{\prime}}=\frac{c_{4}^{3}}{\Delta_{E^{\prime}}}$, $v_{\mathfrak{P}}(j_{E^{\prime}})\equiv-v_{\mathfrak{P}}(\Delta_{E^{\prime}})\pmod{3}$. Hence, $v_{\mathfrak{P}}(j_{E^{\prime}})\equiv 0\text{ or }2\pmod{3}$. This completes the proof of the theorem. ∎
Proof of Theorem 3.1.
Let $(a,b,c)\in\mathcal{O}_{K}^{3}\setminus S$ be a non-trivial primitive solution to the equation (3.1) with exponent $p>V$, where $V=V_{K,A,B,C}$ be the constant as in Theorem 3.5. By Theorem 3.5, there exists an elliptic curve $E^{\prime}/K$ having a non-trivial $3$-torsion point and good reduction away from $S_{K}^{\prime}$. By  (3.2), there exists some $\mathfrak{P}\in S_{K}$ that satisfies $v_{\mathfrak{P}}(\alpha\beta^{-1})=2$.
Now, arguing as in the proof of Theorem 2.5, we find $v_{\mathfrak{P}}(j_{E^{\prime}})\geq 0$ by using $|v_{\mathfrak{P}}(\alpha\beta^{-1})|=2<3v_{\mathfrak{P}}(3)=3$. Recall that $j_{E^{\prime}}=\frac{(\mu+27)(\mu+3)^{3}}{\mu}$, where $\mu=27\alpha\beta^{-1}$. This implies $v_{\mathfrak{P}}(j_{E^{\prime}})\equiv v_{\mathfrak{P}}(\mu+27)-v_{\mathfrak{P}}(\mu)\pmod{3}$. Since $v_{\mathfrak{P}}(\alpha\beta^{-1})=2$, $v_{\mathfrak{P}}(\mu)=5$. Hence $v_{\mathfrak{P}}(j_{E^{\prime}})\equiv 3-5\equiv 1\pmod{3}$. Therefore, $v_{\mathfrak{P}}(j_{E^{\prime}})\geq 0$ and $v_{\mathfrak{P}}(j_{E^{\prime}})\equiv 1\pmod{3}$, which contradicts Theorem 3.5. We are done. ∎
We conclude this section with the following proposition.
Proposition 3.6.
If $A,B,C\in\mathbb{Z}\setminus\{0\}$, then in Theorem 3.1, we can take $S=\emptyset$.
Proof.
In order to prove Theorem 3.1 for $S=\emptyset$, it is enough to prove Theorem 3.2 for $S=\emptyset$. Arguing as in the proof of Theorem 3.2 for $(a,b,c)\in\mathcal{O}_{K}^{3}$, there exists $\mu_{k}\in K$ with $1\leq k\leq n$ such that $E/K$ is modular for all $\mu(E)\notin\{\mu_{1},\mu_{2},...,\mu_{n}\}$, where $j_{E}=3^{3}\frac{(1+\mu(E))(9+\mu(E))^{3}}{\mu(E)^{3}}$ for $\mu(E)=\frac{Bb^{p}}{Aa^{p}}$. WLOG, we can assume $\mu_{1},\mu_{2},...,\mu_{n}\notin\mathbb{Q}^{\ast}$, since elliptic curves over $\mathbb{Q}$ are modular. If $\mu=\mu_{k}$ for some $k\in\{1,2,\ldots,n\}$, then $\left(\frac{b}{a}\right)^{p}=\frac{A\mu_{k}}{B}$. The above equation determines $p$ uniquely, we denote it $p_{k}$. If not, we get $\frac{b}{a}=\pm 1$ and hence $\mu_{k}=\pm\frac{B}{A}\in\mathbb{Q}^{\ast}$, which is a contradiction. Now, argue as in Theorem 3.2 to complete the proof of the proposition. ∎
4. Local criteria of $K$
In this section, we present several local criteria of $K$ which imply Theorem 2.5. First, we look at the case when $K$ is a quadratic field.
Proposition 4.1 (Quadratic).
Let $d\geq 2$ be a normal-$\square$-free integer satisfying $d\equiv 2\pmod{3}$, and let $K=\mathbb{Q}(\sqrt{d})$. Assume $3\nmid h_{K}h_{K(\zeta_{3})}$. If $A,B,C\in\{u3^{r}\ |\ u\in\mathcal{O}_{K}^{\ast},\ r\in\mathbb{Z}_{\geq 0}\}$, then the conclusion of Theorem 2.5 holds over $K$.
Proof.
Since $d\equiv 2\pmod{3}$, $3$ is inert in $K$. Let $S_{K}=\{\mathfrak{P}\}$. The hypotheses on $A,B,C$ imply that $S_{K}^{\prime}=S_{K}$. Now, arguing as in the proof of [Moc22, Theorem 12], we see that the hypothesis of Proposition 2.7 is satisfied. Hence, we are done with the proof of proposition. ∎
Now, we look at the case when $K$ is of odd degree.
Proposition 4.2 (Odd degree).
Let $K$ be a field such that $3\nmid h_{K}h_{K(\zeta_{3})}$ and $n=[K:\mathbb{Q}]$. Suppose
  1. (1)
    $q\geq 5$ be a rational prime with $\gcd(n,q-1)=1$ and totally ramifies in $K$,
  2. (2)
    $3$ is either inert or $3=\mathfrak{P}^{n}$ for some principal ideal $\mathfrak{P}\in P$.
If $A,B,C\in\{u3^{r}\ |\ u\in\mathcal{O}_{K}^{\ast},\ r\in\mathbb{Z}_{\geq 0}\}$, then the conclusion of Theorem 2.5 holds over $K$.
Proof.
Let $\mathfrak{P}\in S_{K}$ be the unique prime ideal lying above $3$. By assumption, $S_{K}$ is principal. The hypotheses on $A,B,C$ imply that $S_{K}^{\prime}=S_{K}$. Now, arguing as in the proof of [Moc22, Theorem 13], we see that the hypothesis of Proposition 2.7 is satisfied. Hence, we are done with the proof of proposition. ∎
Acknowledgments
The authors express their sincere gratitude to Prof. Alain Kraus for his help in understanding the article [Kra90]. The second author thanks ISI Delhi for their generous hospitality during the preparation of this article.